OpenStudy (yanasidlinskiy):

Help please? Which kinematics equation would you use for the following problem: A car traveling at 30m/s begins to gain speed at the rate of 4.0m/s^2. How far will the car travel in 2.0 seconds?

OpenStudy (yanasidlinskiy):

\[v^2=v _{0} ^2+2ad\]

OpenStudy (yanasidlinskiy):

Is it that one?

OpenStudy (phi):

I would expect an equation that shows distance as a function of time.

OpenStudy (yanasidlinskiy):

So.... would this one work then?\[d= 1/2(v+v _{0})t\]

OpenStudy (phi):

I know v= a t d = ½ a t^2 (with initial v0 = 0) d= v t (with initial d0 = 0) so I would use d= ½ a t^2 + v0 t

OpenStudy (phi):

if we put in numbers d = 2 t^2 +30t

OpenStudy (yanasidlinskiy):

So would I subtract 30t?

OpenStudy (phi):

I don't understand the question.

OpenStudy (yanasidlinskiy):

Yea, It's kinda confusing me too. @jigglypuff314 maybe ya can help?

jigglypuff314 (jigglypuff314):

You just needed to indentify the kinematic equation that you want that would be the one with \(v\) (the final/ending velocity) missing \( d = v_0*t + \frac{1}{2}*a*t^2\)